### v1

# Data-Driven Superstabilizing Control under Quadratically-Bounded Errors-in-Variables Noise

Abstract
The Error-in-Variables model of system identification/control involves nontrivial input and measurement corruption of observed data, resulting in generically nonconvex optimization problems. This paper performs full-state-feedback stabilizing control of all discrete-time linear systems that are consistent with observed data for which the input and measurement noise obey quadratic bounds. Instances of such quadratic bounds include elementwise norm bounds (at each time sample), energy bounds (across the entire signal), and chance constraints arising from (sub)gaussian noise. Superstabilizing controllers are generated through the solution of a sum-of-squares hierarchy of semidefinite programs. A theorem of alternatives is employed to eliminate the input and measurement noise process, thus improving tractability. Effectiveness of the scheme is generated on an example system in the chance-constrained set-membership setting where the input and state-measurement noise are i.i.d. normally distributed.
I Introduction
This paper proposes a method to stabilize linear systems corrupted by quadratically bounded Errors-in-Variables (EIV) noise [1]. State and input observations $\mathcal{D}=\{\hat{x}_{t},\hat{u}_{t}\}_{t=0}^{T}$ are collected along a trajectory of an $n$-state $m$-input linear system
$x_{t+1}$ $=Ax_{t}+Bu_{t}$ (1a)
$\hat{x}_{t}$ $=x_{t}+\Delta x_{t},\quad\hat{u}_{t}=u_{t}+\Delta u_{t},$ (1b)
in which the collected data in $\mathcal{D}$ is corrupted by state noise $\Delta x\in\mathbb{R}^{nT}$ and input noise $\Delta u\in\mathbb{R}^{m(T-1)}$. This paper focuses on the setting where the noise processes $(\Delta x,\Delta u)$ satisfy a collection of $N$ convex quadratic constraints as
$\forall\ell$ $\in 1..L:$ $\lVert F_{\ell}\Delta x+G_{\ell}\Delta u\rVert_{2}\leq 1,$ (2)
in which the known constraint matrices $F_{j}$ and $G_{j}$ have compatible dimensions. These quadratic constraints could arise from deterministic knowledge of $(\Delta x,\Delta u)$, or from high-probability chance-constraints imposed on stochastic processes $(\Delta x,\Delta u)$ if a robust description is overly conservative [2, 3]. Our goal is to find a gain matrix $K$ such that the full-state-feedback control policy $u_{t}=Kx_{t}$ can simultaneously stabilize all plants $(A,B)$ that are consistent with the data in $\mathcal{D}$ under the noise description in (2).
This paper follows the framework of set-membership direct Data Driven Control (DDC). In direct DDC, a control policy is formed from the collected data and modeling assumptions without first performing system identification (and synthesizing a controller for the identified system) [4]. Set-membership DDC has three main ingredients: the set of data-consistent plants (given a noise model), the set of commonly stabilized plants by a designed controller, and the certificate of set-containment that the stabilized-set contains the consistent-set [5]. The Matrix S-Lemma can be used to provide proofs of quadratic and worst-case $H_{2}$ or $H_{\infty}$ robust control when the noise model is defined by a matrix ellipsoid (quadratic matrix inequality) [6, 7]. Farkas-based certificates for polytope-in-polytope membership have been used for robust superstabilization [8] and positive-stabilization [9]. Sum of Squares (SOS) certificates of nonnegativity have been employed for stabilization of more general nonlinear systems [10, 11]. We note that other non-set-membership DDC methods include using Virtual Reference Feedback Tuning [12] and Willem’s Fundamental Lemma [13].
Most DDC methods focus solely on process-noise corruption, allowing for the synthesis of controllers through the solution of computationally simple convex programs. This paper continues a line of work in addressing the more challenging setting of EIV superstabilization, proposing a method that can handle the more difficult but realistic EIV noise scenario at the cost of more expensive computational requirements. Our prior work on superstabilization [14, 15] of EIV systems includes full-state-feedback for polytope-bounded noise [16] (e.g. $L_{\infty}$-bounds) and dynamic output feedback for SISO plants [17]. In this work, we will ensure superstabilization under quadratically bounded noise. This will involve developing matrix-SOS expressions defined for (multiple) quadratic constraints in (2). Computational complexity is reduced by eliminating $(\Delta x,\Delta u)$ using a Theorem of Alternatives [18]. The concurrent and independently developed similar work in [19] performs lossless quadratic stabilization in the presence of a single quadratic-matrix-inequality-representable quadratic constraint, with conservatism added under multiple quadratic constraints.
This paper has the following structure: Section II reviews preliminaries including notation, quadratically constrained noise, superstabilization, and matrix-SOS methods. Section III presents infinite-dimensional Linear Programs to perform superstabilization under quadratically-bounded EIV noise, and applies a Theorem of Alternative to eliminate the noise variables $(\Delta x,\Delta u)$ from the linear inequalities. Section IV truncates the infinite-dimensional LPs using the moment-SOS hierarchy of Semidefinite Programs, and tabulates computational complexity of different approaches. Section V incorporates recorded process noise into the quadratically-bounded framework. Section VI demonstrates our method for EIV-tolerant superstabilizing control of example systems. Section VII concludes the paper.
II Preliminaries
BSA
Basic Semialgebraic
DDC
Data Driven Control
EIV
Errors-in-Variables
LP
Linear Program
LMI
Linear Matrix Inequality
PMI
Polynomial Matrix Inequality
PSD
Positive Semidefinite
SDP
Semidefinite Program
SOC
Second-Order Cone
SOS
Sum of Squares
WSOS
Weighted Sum of Squares
II-A Notation
$a..b$
natural numbers between $a$ and $b$ (inclusive)
$\mathbb{R}_{\geq 0}$
nonnegative orthant
$\pi^{x}$
projection operator $(x,y)\rightarrow x$
$A^{\top}$
Transpose of matrix $A$
$A^{+}$
Pseudoinverse of matrix $A$
$\mathbb{L}^{n}$
$n$-dimensional Second-Order Cone (SOC)
$\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}_{\geq 0}\mid t\geq\lVert x\rVert_{2}\}$
$\mathcal{N}(\mathbf{m},\Sigma)$
Normal distribution with mean $\mathbf{m}$
and covariance $\Sigma$
$\mathbb{R}[x]$
polynomials in indeterminate $x$
$\deg(h)$
degree of polynomial $h\in\mathbb{R}[x]$
$\mathbb{S}^{r}[x]$
$r\times r$ polynomial-valued-matrices in $x$
$\Sigma^{r}[x]$
SOS polynomial-valued-matrices in $x$
$\langle A,B\rangle$
Matrix pairing $\textrm{Tr}(A^{\top}B)=\sum_{ij}A_{ij}B_{ij}$
II-B Quadratic Noise Bounds
This subsection briefly highlights instances of quadratic noise bounds in (2).
II-B1 Elementwise Norm Constraints
A deterministic noise bound could impose that $\forall t\in 1..T:\lVert\Delta x_{t}\rVert_{2}\leq\epsilon_{x}$, and $\forall t\in 1..T-1:\lVert\Delta u_{t}\rVert_{2}\leq\epsilon_{u}$. Elementwise norm constraints can arise when $\Delta x$ and $\Delta u$ are i.i.d. normally distributed random variables $(\Delta x_{t}\sim\mathcal{N}(\mathbf{0}_{n},\Sigma_{x}),\Delta u_{t}\sim\mathcal{N}(\mathbf{0}_{m},\Sigma_{u})$. The statistics $\Delta x_{t}^{T}\Sigma_{x}^{-1}\Delta x_{t}$ and $\Delta u_{t}^{T}\Sigma_{u}^{-1}\Delta x_{t}$ are each $\chi^{2}$-distributed with $n$ and $m$ degrees of freedom respectively.
Let $\varepsilon(\delta;f)$ refer to the quantile statistic of a $\chi^{2}$ distribution with $f$ degrees of freedom with respect to a probability $\delta\in[0,1]$ (and random variable $p$):
$\chi^{2}_{f}(p\leq\varepsilon(\delta;f))=1-\delta.$ (3)
For a choice of probabilities $\delta_{x},\delta_{u}\in[0,1]$, the joint probability of $\Delta x,\Delta u$ lying within the set described by
$\forall t$ $\in 1..T:$ $\lVert\Sigma_{x}^{-1/2}\Delta x_{t}\rVert_{2}$ $\leq\sqrt{\varepsilon(\delta_{x};n)}$ (4a)
$\forall t$ $\in 1..T-1:$ $\lVert\Sigma_{u}^{-1/2}\Delta u_{t}\rVert_{2}$ $\leq\sqrt{\varepsilon(\delta_{u};m)}$ (4b)
is $P_{\textrm{joint}}=(1-\delta_{x})^{T}(1-\delta_{u})^{T-1}$. A controller $u=Kx$ certifiably stabilizes all plants consistent with (I) and (II-B1) will be able to stabilize the true system with probability $P_{\textrm{joint}}$. Similar elementwise quadratic chance-constraints arise when $\Delta x,\Delta u$ are drawn from i.i.d. subgaussian distributions [20].
II-B2 Energy Constraints
The standard quadratic expression used in a Linear Quadratic Regulator of
$J=\Delta x_{T}^{\top}Q_{T}\Delta x_{T}+\left(\sum_{t=1}^{T-1}\Delta x_{t}^{\top}Q\Delta x_{t}+\Delta u_{t}^{\top}R\Delta u_{t}\right)$ (5)
is compatible with the structure of (2) if $(Q,R,Q_{T})$ are all Positive Semidefinite (PSD). The mixed structure (for appropriately dimensioned $\tilde{Q},\tilde{R},\tilde{S}$)
$J_{\textrm{mix}}=\begin{bmatrix}\Delta x\\ \Delta u\end{bmatrix}^{T}\begin{bmatrix}\tilde{Q}&\tilde{S}\\ \tilde{S}^{\top}&\tilde{R}\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta u\end{bmatrix}$ (6)
can also be posed as the left-hand-side of the (2) inequality, thus allowing for energy constraints on the noise $(\Delta x,\Delta u)$.
II-C Superstabilizing Control
Let $W\in\mathbb{R}^{f\times n}$ be a matrix with full column rank such that $\{x\mid\lVert Wx\rVert_{\infty}\leq 1\}$ is a compact set. A discrete-time linear system $x_{t+1}=Ax_{t}$ is $W$-superstable if $\lVert Wx\rVert_{\infty}$ is a polyhedral Lyapunov function:
$\lVert W^{+}AW\rVert_{\infty}<1\quad\text{($L_{\infty}$ Operator Norm).}$). (7)
The system is superstable if it is $W$-superstable with $W=I_{n}$. Any superstable system $x_{t+1}=Ax_{t}$ obeys the the decay bound of $\lVert x_{t}\rVert_{\infty}\leq\lVert A\rVert_{\infty}^{t}\lVert x_{0}\rVert_{\infty}$. This decay bound generalizes to $W$-superstability as in $\lVert Wx_{t}\rVert_{\infty}\leq\lVert W^{+}AW\rVert_{\infty}^{t}\lVert Wx_{0}\rVert_{\infty}$. $W$-superstabilization of the system $x_{t+1}=Ax_{t}+Bu_{t}$ proceeds by choosing a gain $K\in\mathbb{R}^{m\times n}$ with $u_{t}=Kx_{t}$ such that $\lVert A+BK\rVert_{\infty}<1$. The superstabilization problem of minimizing the decay bound is a finite-dimensional LP:
$\lambda^{*}=$ $\inf_{M,K}\quad\lambda:\qquad\textstyle\sum_{j=1}^{f}M_{ij}<\lambda$ $\forall i\in 1..f$ (8a)
$\textstyle\lvert[W^{+}(A+BK)W]_{ij}\rvert\leq M_{ij}$ $\forall i,j\in 1..f.$ (8b)
$M\in\mathbb{R}^{n\times n},\ K\in\mathbb{R}^{m\times n}.$ (8c)
II-D Sum-of-Squares Matrices Background
This paper will formulate worst-case-superstabilization problems as infinite-dimensional LPs, which in turn will be truncated using the moment-SOS hierarchy [21]. For any symmetric polynomial-valued-matrix $p\in\mathbb{S}^{r}[x]$ of size $r\times r$ with indeterminate $x\in\mathbb{R}^{n}$, the degree of $p$ is the maximum polynomial degree of any one of its entries $(\deg p=\max_{ij}\deg p_{ij}).$ A sufficient condition for $p(x)\succeq 0$ over $\mathbb{R}^{n}$ is if $p(x)$ is an SOS-matrix: there exists a polynomial vector $v\in(\mathbb{R}[x])^{c}$ and a PSD Gram matrix $Q\in\mathbb{S}^{rc\times rc}$ such that (Lemma 1 of [22])
$p(x)=(v(x)\otimes I_{r})^{\top}Q(v(x)\otimes I_{r}).$ (9)
The set of SOS matrices with representation in (9) is $\Sigma^{r}[x]$, and the subset of SOS matrices with maximal degree $\leq 2k$ is $\Sigma^{r}[x]_{\leq 2k}$. A constraint region defined by a locus of PSD containments can be constructed from $\{g_{j}\in\mathbb{S}^{r_{j}}[x]\}_{j=1}^{N_{c}}$ as
$\mathbb{K}=\{x\in\mathbb{R}^{n}\mid\forall j\in 1..N_{c}:\ g_{j}(x)\succeq 0\}.$ (10)
The matrix $p\in\mathbb{S}^{r}[x]$ satisfies \@iaciPMI Polynomial Matrix Inequality (PMI) over the region in (10) if $\forall x\in\mathbb{K}:\ p(x)\succeq 0.$ A sufficient condition for this PMI to hold is if there exists SOS-matrices $\{\sigma_{j}(x)\}_{j=0}^{N_{c}}$ such that
$p(x)$ $=\sigma_{0}(x)+\textstyle\sum_{i=j}^{N_{c}}\langle g_{j}(x),\sigma_{j}(x)\rangle$ (11a)
$\sigma_{0}\in\Sigma^{r}[x],\ \forall j\in 1..N_{c}:\ \sigma_{j}\in\Sigma^{r_{j}}[x],\epsilon>0.$ (11b)
The set of matrices in $\mathbb{S}^{r}[x]$ possessing a representation as in (11a) (existence of $\{\sigma_{j}\}$) will be written as the Weighted Sum of Squares (WSOS) set $\Sigma^{r}[\mathbb{K}]$. The degree-$2k$-bounded WSOS cone $\Sigma^{r}[\mathbb{K}]_{\leq 2k}$ is the set of matrices with representation in (II-D) such that $\deg\sigma_{0}\leq 2k$ and $\forall j\in 1..N_{c}:\ \deg\langle g_{j}(x),\sigma_{j}(x)\rangle\leq 2k$. The representation of $\mathbb{K}$ by polynomial matrices in (10) is Archimedean if there exists an $R>0$ such that the scalar polynomial $p_{R}(x)=R-\lVert x\rVert_{2}^{2}$ satisfies $p_{R}\in\Sigma^{r}[\mathbb{K}]$. If the representation in (10) is Archimedean, then every $p\in\mathbb{S}^{r}[x]$ such that $p(x)$ is Positive Definite over $K$ satisfies $(p(x)-\epsilon I_{r})\in\Sigma^{r}[K]$ for some $\epsilon>0$ (Corollary 1 of [22]). Testing membership of $p\in\Sigma^{r}[x]_{\leq 2k}$ through (II-D) can be accomplished by solving \@iaciLMI Linear Matrix Inequality (LMI) in the Gram matrices for $\{\sigma_{j}\}$ under coefficient-matching equality constraints. A common choice of polynomial vector $v(x)$ used to represent the SOS matrices in (9) is the vector of all $\binom{n+k}{k}$ monomials of degree $\leq k$. The maximal-size PSD constraint involved in (II-D) under the monomial choice for $v(x)$ is $r\binom{n+k}{k}$.
III Quadratically-Bounded Linear Programs
This section presents an infinite-dimensional LP which $W$-superstabilizes the class of data-consistent plants.
III-A Consistency Set
Define $h^{0}$ as the following residual:
$h_{t}^{0}$ $=\hat{x}_{t+1}-A\hat{x}_{t}-B\hat{u}_{t}$ $\forall t\in 1..T-1.$ (12)
The quantity $h^{0}$ will be equal to zero at all $t\in 1..T-1$ if there is no noise present in the system.
The joint consistency set $\bar{\mathcal{P}}(A,B,\Delta x,\Delta u)\subset\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times m}\times\mathbb{R}^{nT}\times\mathbb{R}^{m(T-1)}$ of plants and noise values consistent with the dynamics (1a) and noise properties (2) is
$\bar{\mathcal{P}}:=\ \begin{Bmatrix}[l]\Delta x_{t+1}=A\Delta x_{t}+B\Delta u_{t}+h_{t}^{0}&\forall t\in 1..T-1\\ \lVert F_{\ell}\Delta x+G_{\ell}\Delta u\rVert_{2}\leq 1&\forall\ell\in 1..L\\ \end{Bmatrix},$ (15)
The formulation (15) involves bilinear terms $A\Delta x$ and $B\Delta u$, which could lead to nonconvex or possibly disconnected sets $\bar{\mathcal{P}}$ [23]. By defining a matrix $\Xi_{A}\in\mathbb{R}^{n(T-1)\times nT}$ with
$\Xi$ $=\begin{bmatrix}A&-I&\mathbf{0}&\ldots\\ \mathbf{0}&A&-I&\ldots\\ \vdots&\vdots&\vdots&\ddots\end{bmatrix}$ (19)
$=[(I_{T-1}\otimes A),\mathbf{0}_{n(T-1)\times n}]+[\mathbf{0}_{n(T-1)\times n},-I_{n(T-1)}],$
and defining $n_{\ell}$ to be the column dimension of $F_{\ell}$ and $G_{\ell}$ for each $\ell$ ($F_{\ell}\in\mathbb{R}^{n_{\ell}\times nT},\ G_{\ell}\in\mathbb{R}^{n_{\ell}\times m(T-1)}$, the consistency set in (15) can be equivalently expressed as
$\bar{\mathcal{P}}:=\ \begin{Bmatrix}[l]\Xi\Delta x+(I_{T-1}\otimes B)\Delta u+h_{0}=0\\ (F_{\ell}\Delta x+G_{\ell}\Delta u,1)\in\mathbb{L}^{n_{\ell}}&\forall\ell\in 1..L\\ \end{Bmatrix}.$ (22)
The consistency set $\mathcal{P}(A,B)\subset\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times m}$ of plants compatible with the data in $\mathcal{D}$ is the following projection,
$\mathcal{P}(A,B)=\pi^{A,B}\bar{\mathcal{P}}(A,B,\Delta x,\Delta u),$ (23)
in which there exists bounded noise $(\Delta x,\Delta u)$ following (2) such that $\mathcal{D}$ could have been observed. Our task is as follows:
Problem 1
Find a matrix $K\in\mathbb{R}^{m\times n}$ such that full-state-feedback controller $u_{t}=Kx_{t}$ ensures that $A+BK$ is $W$-superstable for all $(A,B)\in\mathcal{P}$.
III-B Full Linear Program
Let $\eta>0$ be chosen as a margin to ensure stability.
Theorem III.1
$W$-Superstability through a common $K$ (for Problem 1) can be ensured by solving the following infinite-dimensional LP in terms of a matrix $K\in\mathbb{R}^{m\times n}$ and a matrix-valued function $M(A,B):\mathcal{P}\rightarrow\mathbb{R}^{f\times f}$ (Equation (12) of [24]):
$\operatorname*{find}_{M,K}$ $\forall(A,B,\Delta x,\Delta u)\in\bar{\mathcal{P}}(A,B,\Delta x,\Delta u):$ (24a)
$\quad\forall i=1..f:$ (24b)
$\qquad 1-\eta-\textstyle\sum_{j=1}^{n}M_{ij}(A,B)\geq 0$
$\quad\forall i=1..f,\ j=1..f:$ (24c)
$\qquad M_{ij}(A,B)-(W^{+}\left(A+BK\right)W)_{ij}\geq 0$
$\qquad M_{ij}(A,B)+(W^{+}\left(A+BK\right)W)_{ij}\geq 0$
$M(A,B):\mathcal{P}\rightarrow\mathbb{R}^{f\times f},\ K\in\mathbb{R}^{m\times n}.$ (24d)
Proof:
The $W$-superstabilizing linear system in (II-C) is fulfilled for each $(A,B)\in\mathcal{P}$, as expressed by the $\forall(A,B,\Delta x,\Delta u)\in\bar{\mathcal{P}}$ quantification in (24a). ∎
Theorem III.2
Under the assumption that $\bar{\mathcal{P}}$ is compact, the function $M$ can be selected to be continuous (Lemma 3.1 of [24]) and even polynomial (Lemma 3.2 of [24])
Proof:
This follows from Lemmas 3.1 (continuity) and 3.2 (polynomial approximability) of [24]. The same type of linear system structure for superstabilization in (24b)-(24d) is used (with $W=I_{n}$), but the set $\bar{\mathcal{P}}$ in [24] is defined by polytopic-bounded noise rather than quadratically-bounded noise (15). Proof details are omitted for space. ∎
III-C Robustified Linear Program
Problem (III.1) involves $2f^{2}+f$ infinite-dimensional LP, each posed over the $n(n+T)+m(n+T-1)$ variables $(A,B,\Delta x,\Delta u)$. Reducing the number of variables involved in any quantification will simplify resultant SOS truncations and increase computational tractability.
This subsection will use a Theorem of Alternatives in order to eliminate the noise variables $(\Delta x,\Delta u)$ from Program (III.1). The maximum number of variables appearing in any infinite-dimensional linear inequality constraint will subsequently drop from $n(n+T)+m(n+T-1)$ down to $n(n+m)$. Let $q(A,B):\mathcal{P}\rightarrow\mathbb{R}$ be a function that is independent of $(\Delta x,\Delta u)$, such as any one of the left-hand-side elements for constraints (24b)-(24c). The following pair of problems are strong alternatives (exactly one is feasible):
$\forall(A,B,\Delta x,\Delta u)\in\bar{\mathcal{P}}:$ $q(A,B)$ $\geq 0$ (25a)
$\exists(A,B,\Delta x,\Delta u)\in\bar{\mathcal{P}}:$ $q(A,B)$ $<0,$ (25b)
because if $q\geq 0$ for all $(A,B,\Delta x,\Delta u)\in\bar{\mathcal{P}}$ (25a), then there cannot be a point in $\bar{\mathcal{P}}$ where $q(A,B,\Delta x,\Delta u)<0$ (25b). We can define the following dual variable functions
$\mu:$ $\mathcal{P}\rightarrow\mathbb{R}^{n(T-1)}$ (26a)
$(s_{\ell},\tau_{\ell}):$ $\mathcal{P}\rightarrow\mathbb{L}^{n_{\ell}}$ $\forall\ell\in 1..L.$ (26b)
The $(A,B)$ dependence of $(\mu,s,\tau)$ will be omitted to simplify notation. A weighted sum $\Phi(A,B,\Delta x,\Delta u;\mu,s,\tau)$ may be constructed from $q$ and $(\mu,s,\tau)$ from (III-C):
$\Phi$ $=-q+\mu^{\top}(\Xi\Delta x+(I_{T-1}\otimes B)\Delta u+h^{0})$
$\textstyle+\sum_{\ell=1}^{L}\left(s_{\ell}^{\top}(F_{\ell}\Delta x+G_{\ell}\Delta u)-\tau_{\ell}\right).$ (27)
The terms in $\Phi$ may be rearranged into
$\Phi$ $=\textstyle\left(-q+\mu^{\top}h^{0}-\sum_{\ell=1}^{L}\tau_{\ell}\right)$
$\textstyle+\left(\Xi^{\top}\mu+\sum_{\ell=1}^{L}F_{\ell}^{\top}s_{\ell}\right)^{\top}\Delta x$ (28)
$+\textstyle\left((I_{T-1}\otimes B^{\top})\mu+\sum_{\ell=1}^{L}G_{\ell}^{\top}s_{\ell}\right)^{\top}\Delta u.$
Expressing the $(\Delta x,\Delta u)$-constant terms of $\Phi$ as $Q$ with
$Q=-q+\mu^{\top}h^{0}-\sum_{\ell=1}^{L}\tau_{\ell},$ (29)
the supremal value of $\Phi$ w.r.t. $\Delta x,\Delta u$ has a value of
$\sup_{\Delta x,\Delta u}\Phi=\begin{cases}Q&\textrm{if}\ \Xi^{\top}\mu+\sum_{\ell=1}^{L}F_{\ell}^{\top}s_{\ell}\\ &\textrm{if}\ (I_{T-1}\otimes B^{\top})\mu+\sum_{\ell=1}^{L}G_{\ell}^{\top}s_{\ell}\\ \infty&\textrm{otherwise}.\end{cases}$ (30)
Theorem III.3
Problem (25a) will have the same feasibility (or infeasibility) status as the following program:
$\operatorname*{find}_{\mu,s,\tau}$ $\forall(A,B)\in\mathcal{P}:$ (31a)
$\qquad q+\textstyle\sum_{\ell=1}^{L}\tau_{\ell}-\mu^{\top}h^{0}\geq 0$ (31b)
$\textstyle\qquad\Xi^{\top}\mu+\sum_{\ell=1}^{L}F_{\ell}^{\top}s_{\ell}=0$ (31c)
$\qquad\textstyle(I_{T-1}\otimes B^{\top})\mu+\sum_{\ell=1}^{L}G_{\ell}^{\top}s_{\ell}=0$ (31d)
$(\mu,s,\tau)$ from (III-C). (31e)
Proof:
This relationship holds using the convex-duality based Theorem of Alternatives from Section 5.8 of [18], given that all description expressions in $q$ and $\bar{\mathcal{P}}$ are affine in the uncertain terms $(\Delta x,\Delta u)$. ∎
Theorem III.4
If $\mathcal{P}$ is compact, then the multipliers $(\mu,s,\tau)$ can be chosen to be polynomial functions of $(A,B)$.
Proof:
The proof is omitted due to page limitations, but follows from Theorem 3.3 of [25] (generalizing Theorems 4.4 and 4.5 of [16] to the conically constrained case). ∎
In the case of probabilistic noise set from (II-B1) with $L=2T-1$, the multipliers $(s,\tau)$ can be partitioned into $\forall t\in 1..T:\ (s^{x}_{t},\tau^{x}_{t})$ for (4a) and $\forall t\in 1..T-1:\ (s^{u}_{t},\tau^{u}_{t})$ for (4b). The certificate (III.3) can then be expressed as
$\operatorname*{find}_{\mu,s,\tau}$ $\forall(A,B)\in\mathcal{P}:$ (32a)
$\qquad q+\textstyle\sum_{\ell=1}^{T}\sqrt{\varepsilon(\delta_{x};n)}\tau^{x}_{\ell}+\textstyle\sum_{\ell=1}^{T-1}\sqrt{\varepsilon(\delta_{u};m)}\tau^{u}_{\ell}$
$\qquad\qquad-\mu^{\top}h^{0}\geq 0$ (32b)
$\qquad\Sigma^{-1/2}_{x}s_{1}^{x}=-A^{\top}\mu_{1}$ (32c)
$\forall t=2..T-1:\ \Sigma^{-1/2}_{x}s_{t}^{x}=\mu_{t-1}-A^{\top}\mu_{t}$ (32d)
$\qquad\Sigma^{-1/2}_{x}s_{T}^{x}=\mu_{T-1}$ (32e)
$\forall t\in 1..T-1:\ \Sigma_{u}^{-1/2}s^{u}_{t}=-B^{\top}\mu_{t}$ (32f)
$(\mu,s,\tau)$ from (III-C). (32g)
IV Truncated Sum-of-Squares Programs
This section uses the moment-SOS hierarchy of SDPs to discretize the infinite-dimensional LP in (III.1) into finite-dimensional convex optimization problems that are more amenable to computation. This discretization will be performed by SOS-matrix truncations.
IV-A Quadratically-Bounded Truncations
Program (III.1) has $2f^{2}+f$ infinite-dimensional LP constraints posed over $(A,B,\Delta x,\Delta u)$. This subsection applies a degree $2k$ SOS tightening to the constraints in program (III.1). In each case, the matrix $M$ is restricted to a polynomial $M\in(\mathbb{R}[A,B])^{n\times n}_{\leq 2k}.$ The remainder of this section will analyze the computational complexity of an SOS tightening on a single infinite-dimensional constraint from (III.1) (represented as $q(A,B)\geq 0$ from (25a)). Complexity will be compared according to the PSD Gram matrix of maximal size.
IV-A1 Full Program
The full program applies a scalar WSOS constraint $q\in\Sigma^{1}[\bar{\mathcal{P}}]$ posed over variables $(A,B,\Delta x,\Delta u)$. The size of the maximal Gram matrix for each (25a) restriction is
$p_{F}=\binom{n(n+T)+m(n+T-1)+k}{k}.$ (33)
IV-A2 Alternatives (Dense)
The SOC constraint in (26b) can also be expressed by a PSD constraint [26],
$(s_{\ell},\tau_{\ell})\in\mathbb{L}^{n_{\ell}}$ $\Leftrightarrow\begin{bmatrix}\tau_{\ell}I_{n_{\ell}}&s_{\ell}\\ s_{\ell}^{\top}&\tau_{\ell}\end{bmatrix}\in\mathbb{S}^{n_{\ell}+1}_{+}.$ (36)
In order to prove convergence of the Alternative truncations as the degree $k\rightarrow\infty$, we must assume that there exists a known Archimedean set $\Pi(A,B)$ such that $\Pi\supseteq\mathcal{P}$. Such a $\Pi$ might be known from norm or Lipschitz bounds on $(A,B)$, or other similar knowledge on reasonable plant behavior. If $\Pi$ is a-priori unknown, then we will use the WSOS symbol $\Sigma^{r}[\Pi]$ to refer to the SOS set $\Sigma^{r}[x]$. At the degree-$k$ truncation, the multipliers from (III-C) can be chosen:
$\mu$ $\in(\mathbb{R}[\Pi]_{\leq 2k-1})^{nT}$ (37a)
$s_{\ell}$ $\in(\mathbb{R}[\Pi]_{\leq 2k})^{n_{\ell}}$ $\forall\ell\in 1..L$ (37b)
$\tau_{\ell}$ $\in\mathbb{R}[\Pi]_{\leq 2k}$ $\forall\ell\in 1..L.$ (37c)
These multipliers are required to satisfy:
$\begin{bmatrix}\tau_{\ell}(A,B)I_{n}&s_{\ell}(A,B)\\ s_{\ell}(A,B)^{\top}&\tau_{\ell}(A,B)\end{bmatrix}\in\Sigma^{n_{\ell}+1}[\Pi]_{\leq 2k}$ $\forall\ell\in 1..L.$ (37f)
The matrix WSOS constraint in (37f) has \@iaciPSD PSD Gram matrix of maximal size
$p_{A}^{\ell}=(n^{\ell}+1)\binom{n(n+m)+k}{k}.$ (38)
The ‘dense’ nomenclature for this approach will refer to the imposition that (37f) is WSOS over a matrix of size $n_{\ell}+1$ for each $\ell=1..L.$
IV-A3 Alternatives (Sparse)
The SOC constraint in (36) can be decomposed into $2\times 2$ blocks as in [26]:
$(s_{\ell},\tau_{\ell})\in\mathbb{L}^{n_{\ell}}$ $\Leftrightarrow\exists z_{i\ell}:\ \begin{bmatrix}\tau_{\ell}&s_{i\ell}\\ s_{i\ell}&z_{i\ell}\end{bmatrix}\in\mathbb{S}^{2}_{+},\quad\tau_{\ell}=\textstyle\sum_{i=1}^{n}z_{i\ell}.$ (41)
The multipliers in (III-C) can be restricted to
$\mu$ $\in(\mathbb{R}[\Pi]_{\leq 2k-1})^{nT}$ (42a)
$s_{\ell}$ $\in(\mathbb{R}[\Pi]_{\leq 2k})^{n_{\ell}}$ $\forall\ell\in 1..L$ (42b)
$z_{\ell}$ $\in(\mathbb{R}[\Pi]_{\leq 2k})^{n_{\ell}}$ $\forall\ell\in 1..L$ (42c)
$\tau_{\ell}$ $=\textstyle\sum_{i=1}^{n_{\ell}}z_{i\ell}$ $\forall\ell\in 1..L$ (42d)
subject to the constraint
$\begin{bmatrix}\textstyle\tau_{\ell}(A,B)&s_{i\ell}(A,B)\\ s_{i\ell}(A,B)&z_{i\ell}(A,B)\end{bmatrix}\in\Sigma_{\leq 2k}^{2}[\Pi]$ $\forall\ell=1..L.$ (42g)
The Gram matrices from (42g) have a reduced size of
$p_{B}=2\binom{n(n+m)+k}{k}.$ (43)
Remark 1
The cone constraints (37f) and (42g) are not necessarily equivalent at finite degree $2k$, although they will describe the same set as $k\rightarrow\infty$ given that $(s,\tau)$ are optimization variables. Refer to [27] for more details on the relationship between matrix SOC cone representations.
IV-B Computational Complexity
Computational complexity of the Full, Alternatives (Dense), and Alternatives (Decomposed) schemes will be judged comparing the sizes and multiplicities of the largest PSD constraint in any $\bar{\mathcal{P}}$-nonnegativity constraint. As a reminder, the extended superstabilizing program in (III.1) has $2n^{2}+n$ such $\bar{\mathcal{P}}$-nonnegativity linear inequality constraints.
For each $\bar{\mathcal{P}}$-nonnegativity constraint, the Full program requires 1 block of maximal size $p_{F}$ from (33). The Alternatives programs require $L+1$ PSD blocks each, in which the Dense program has block sizes of $p_{A}^{\ell}$ from (38) for $\ell=1..L$ (and a scalar block of size $\binom{n(n+m)+k}{k}$). The sparse program has all $L$ blocks of size $p_{B}$ from (43) along with a similar scalar block of size $\binom{n(n+m)+k}{k}$. Table I reports the size of the largest PSD matrix constraint for the three approaches under $n=2,m=2,k=2,T=12$. The considered quadratically-bounding constraints all have $L=2$, such as in the elementwise-norm $\lVert\Delta x_{t}\rVert_{2}\leq\epsilon_{x}$ and $\lVert\Delta u_{t}\rVert_{2}\leq\epsilon_{u}$ from Section II-B1. We first note that the per-iteration complexity of solving an SDP using an interior point method (with PSD size $N$ having $M$ affine constraints) is $O(N^{3}M+N^{2}M^{2})$ [28]. While the specifics of parameter scaling for nonsymmetric interior point SOS solvers are different (lacking the affine constraint embedding) [29], the maximal PSD size remains a valuable way to compare predicted performance. In the context of Table I, the size of $N=p_{F}=1540$ is intractably large for current interior-point methods. The ‘Count’ scaling for Dense and Sparse causes computational complexity to grow linearly, while the increasing ‘Size’ parameter causes polynomial growth in scaling.
TABLE I: Size of PSD Variables for $\bar{\mathcal{P}}$-nonnegativity
Size Count
Full $p_{F}=1540$ 1
Dense $p_{A}^{\ell}=135$ 46
Sparse $p_{B}=90$ 46
V Included Process Noise
Process noise can be added to the quadratically-bounded EIV framework developed in this paper. Consider structured process noise $\{w_{t}\}$ corrupting dynamics in (1a), leading to the relation (with known $E\in\mathbb{R}^{n\times n_{w}}$)
$x_{t+1}=Ax_{t}+Bu_{t}+Ew_{t}.$ (44)
The consistency set of plants $\bar{\mathcal{P}}^{w}(A,B,\Delta x,\Delta u,w)$ compatible with data in $\mathcal{D}$ is
$\bar{\mathcal{P}}^{w}:=\begin{Bmatrix}[l]\Xi\Delta x+(I_{T-1}\otimes B)\Delta u+(I_{T-1}\otimes E)w=-h^{0}\\ (F_{\ell}\Delta x+G_{\ell}\Delta u+H_{\ell}w,1)\in\mathbb{L}^{n_{\ell}}\qquad\forall\ell\in 1..L\\ \end{Bmatrix},$ (47)
in which $\{H_{\ell}\}$ are matrices reflecting the quadratic constraints imposed on $w$. The process-noise version of the constraint (25a) is
$\forall(A,B,\Delta x,\Delta u,w)\in\bar{\mathcal{P}}^{w}:$ $q(A,B)\geq 0.$ (48)
A degree-$2k$ SOS truncation of (48) can proceed directly by $q(A,B)\in\Sigma^{1}[\bar{\mathcal{P}}^{w}]_{\leq 2k},$ yielding Gram matrices of size $\binom{n(n+T)+m(n+T-1)+n_{w}+k}{k}$. The Theorem of Alternatives process eliminates the noise variables $(\Delta x,\Delta u,w)$ altogether. The expression of (48) adds the following entry to the constraints in (III.3):
$(I_{T-1}\otimes E^{\top})\mu+\textstyle\sum_{\ell=1}^{L}H_{\ell}^{\top}s_{\ell}=0.$ (49)
Just like in the case without process noise, the infinite-dimensional linear constraints depend solely on $(A,B)$ after robustifying against noise, and the size of the maximal Gram matrices are found in (38) and (43).
VI Numerical Example
Code to generate the following experiment is available at https://github.com/Jarmill/eiv_quad. The SDPs deriving from the SOS programs were synthesized through JuMP [30] and solved using Mosek 10.1 [31].
This example involves a 2-state 2-input ground-truth plant:
$A_{\star}=\begin{bmatrix}0.6863&0.3968\\ 0.3456&1.0388\end{bmatrix},\ B_{\star}=\begin{bmatrix}0.4170&0.0001\\ 0.7203&0.3023\end{bmatrix}.$ (54)
Data $\mathcal{D}$ is collected from an execution of (54) over a time-horizon of $T=14$. In the data collection, the EIV noise $\Delta x,\Delta u$ are i.i.d. distributed according to $\Delta x_{t}\in\mathcal{N}(\mathbf{0},0.04^{2}I_{2}),\Delta u_{t}\in\mathcal{N}(\mathbf{0},0.03^{2}I_{2})$. It is desired to create a superstabilizing $(W=I_{2})$ controller that will succeed in regulating the ground-truth system with joint probability $P{{}_{\textrm{joint}}}=95\%.$ The per-noise probability is chosen as $\delta_{x}=\delta_{u}=(0.95)^{1/(2T-1)}=0.9981$. The $(\delta_{x},\delta_{u})$ chance constraint is modeled as $\lVert\Delta x_{t}\rVert_{2}\leq 0.04\epsilon(\delta_{x};2)=0.1416$ and $\lVert\Delta u_{t}\rVert_{2}\leq 0.03\epsilon(\delta_{u};2)=0.1062$ for each $t$.
Superstabilizing control is performed under a decay-bound objective $\lambda^{*}=\min_{\lambda,K}\lambda\geq\lVert A+BK\rVert_{\infty}\ \forall(A,B)\in\mathcal{P}$. The $k=1$ truncation of the Alternatives certification (IV-A2) (dense, $p_{A}=27$) returns the following policy:
$K_{\textrm{dense}}$ $=\begin{bmatrix}-1.1559&-0.1193\\ 1.1333&-0.9109\end{bmatrix}$ $\lambda^{*}_{\textrm{dense}}=0.9907.$ (57)
Worst-case superstabilization is certified by the value $\lambda^{*}<1$. The ground-truth system (54) has a decay-bound of $\lambda_{\star}=\lVert A_{\star}+B_{\star}K_{\textrm{dense}}\rVert_{\infty}=0.8219$ with respect to the controller $K$ in (57). The Full program ($p_{F}=62$) and Sparse ($p_{B}=18$) fail to perform superstabilization at $k=1$ because the returned bound satisfies $\lambda^{*}_{\textrm{full}}=3.1749\times 10^{3},\lambda^{*}_{\textrm{sparse}}=2.4173\geq 1$. Attempting execution of the $k=2$ tightening for the Full program ($p_{F}=\binom{64}{2}=2016$) results in out-of-memory errors in Mosek. For this specific example, the algorithms from Theorems 1 and 2 of [19] both fail to find a common quadratically stabilizing controller on (matrix-ellipsoid-based over-approximations of) $\mathcal{P}$.
VII Conclusion
This paper presented a solution approach for data-driven superstabilization in the quadratically-bounded EIV setting. The $W$-superstabilization problem was formulated as an infinite-dimensional linear program, and was discretized using SOS-matrix methods. A Theorem of Alternative was used to eliminate the $(\Delta x,\Delta u)$ noise terms, resulting in matrix SOS constraints with lower computational complexity. Future work involves decreasing the runtime involved in posing solutions to EIV superstabilization, such as by investigating other approaches for discretization of the infinite-dimensional LPs. Other aspects include incorporating streaming data for online superstabilization, which could include integrating the EIV method into model predictive control.